# Multipole Attention for Efficient Long Context Reasoning

### v1

@@ Abstract
Large Reasoning Models (LRMs) have shown promising accuracy improvements on complex problem-solving tasks. While these models have attained high accuracy by leveraging additional computation at test time, they need to generate long chain-of-thought reasoning in order to think before answering, which requires generating thousands of tokens. While sparse attention methods can help reduce the KV cache pressure induced by this long autoregressive reasoning, these methods can introduce errors which disrupt the reasoning process. Additionally, prior methods often pre-process the input to make it easier to identify the important prompt tokens when computing attention during generation, and this pre-processing is challenging to perform online for newly generated reasoning tokens. Our work addresses these challenges by introducing Multipole Attention, which accelerates autoregressive reasoning by only computing exact attention for the most important tokens, while maintaining approximate representations for the remaining tokens. Our method first performs clustering to group together semantically similar key vectors, and then uses the cluster centroids both to identify important key vectors and to approximate the remaining key vectors in order to retain high accuracy. Additionally, we design a fast cluster update process to quickly re-cluster the input and previously generated tokens, thereby allowing for accelerating attention to the previous output tokens. We evaluate our method using emerging LRMs such as Qwen-8B and Deepseek-R1-Distil-Qwen2.5-14B, demonstrating that our approach can maintain accuracy on complex reasoning tasks even with aggressive attention sparsity settings. We also provide kernel implementations to demonstrate the practical efficiency gains from our method, achieving up to 4.5 speedup for attention in long-context reasoning applications. Our code is available at https://github.com/SqueezeAILab/MultipoleAttention.
@@ 1 Introduction
Test-time compute has emerged as a new dimension for scaling up Large Language Model (LLM) performance Snell et al. (2024). By scaling the amount of computation used at test time, we can allow LLMs to think longer on harder problems in order to attain higher accuracy. Large Reasoning Models (LRMs), which are LLMs post-trained to have strong reasoning capabilities, typically reason by generating long chain-of-thought, where they produce a step-by-step chain to help guide themselves to the correct solution Wei et al. (2022); Guo et al. (2025); Qwen (2025). This has been shown to provide substantial accuracy gains in domains such as math and coding which require problem solving capabilities Guo et al. (2025); Qwen (2025).
However, the long autoregressive generation required for reasoning also brings substantial inference efficiency challenges. Notably, the reasoning process leads to large Key/Value (KV) cache memory footprint Hooper et al. (2024b). One common method to reduce the KV cache footprint is through sparse attention methods that only load important KV cache tokens at each generation step Tang et al. (2024); Hooper et al. (2024a). However, while these methods are promising for reducing the KV cache memory bottleneck, they lead to unacceptable accuracy loss. Prior work has also shown that problem solving tasks which require complex reasoning experience greater accuracy degradation when pruning the KV cache Liu et al. (2025). Additionally, existing sparse attention methods often pre-process the KV cache from the prompt to make it easier to retrieve important prompt tokens during generation, and this pre-processing is challenging to perform online during generation for the previously generated reasoning tokens Hooper et al. (2024a).
To address these challenges, we build on top of Hooper et al. (2024a) and introduce Multipole Attention, which leverages sparse attention in order to only load the important KV cache tokens, while still approximating attention to the rest of the context to maintain high accuracy. Figure 1 provides an intuitive visualization of our algorithm, which first clusters the keys and computes a representative centroid for each cluster. When computing attention, our method compares the current query with the centroids to determine the importance of key tokens. For the highest scoring keys, we compute exact attention. For the remaining keys, we leverage the cluster centroid to approximate the attention score for all keys in that cluster, thereby retaining contextual information from the full KV cache. We also present a hierarchical generalization of our method which leverages progressively coarser grained centroids as we get further away from the query. To accelerate attention to the previously generated reasoning tokens, we also design a fast cluster update strategy that can be applied during generation, thereby allowing our method to accelerate attention to the previous output tokens.
@@
Multipole Approximation: Our algorithm clusters keys based on semantic similarity and represents all keys within a cluster with a representative key centroid (as in Hooper et al. (2024a)). When computing attention, we then compare the current query with the key centroids to identify which keys are important for the current query; we only compute exact attention for these important keys, and we approximate the attention to the remaining keys using the attention scores with the cluster centroids. This allows our algorithm to maintain contextually relevant information from the full sequence, even with low KV cache budget. We also extend this to a hierarchical multipole approximation to improve the efficiency of the centroid comparison.
@@ 2.1 Reasoning Models
Reasoning models have emerged as a new paradigm for solving complex problem-solving tasks Guo et al. (2025); Qwen (2025). These models are trained to generate a chain-of-thought reasoning trajectory, which contains step-by-step reasoning to help guide the model to the correct answer Wei et al. (2022). The DeepSeek-R1 model Guo et al. (2025) pioneered a new reinforcement learning-based methodology for training reasoning models for problem solving tasks. Their model release also included smaller open-source models (distilled from their base model) based on the Llama3 Grattafiori et al. (2024) and Qwen2.5 model series Yang et al. (2024). There have also been several powerful reasoning models released with agentic tool-calling capabilites such as the QwQ-32B model Qwen (2025), the Qwen3 model series Yang et al. (2024), and the Llama-Nemotron model series Bercovich et al. (2025).
@@ 2.2 Long Context Inference
Long-context length LLMs, which can support context lengths greater than 100K tokens, have seen widespread use for applications such as document question answering, summarization, literature review, and processing multi-turn conversations. There have been several closed-source models which support long context capabilities Achiam et al. (2023); Anthropic (2023), including Gemini Google (2023) which supports greater than 1 million context length. Newer open-source model families such as Llama3 Meta (2024) and Qwen2.5 Yang et al. (2024) also support greater than 100K context length. Another line of work has focused on extending the context windows of pretrained models beyond the original supported context length through scaling or adjusting the positional embeddings Chen et al. (2023); Peng et al. (2023).
@@ 2.3 KV Cache Compression
KV cache compression has emerged as a key algorithmic tool for enabling long context length inference. For long context inference scenarios, the KV cache becomes the main memory bottleneck, making KV cache compression crucial for efficient inference Hooper et al. (2024b). Common approaches for compressing the KV cache include quantization Hooper et al. (2024b); Liu et al. (2024b), which aims to compress the representation for each token, and sparsification Tang et al. (2024); Hooper et al. (2024a); Zhang et al. (2024b), which aims to only load important KV cache tokens.
@@
One previous approach for inducing sparsity to reduce KV cache size during inference is KV cache eviction, where less important tokens are evicted during the generation process to reduce memory consumption. Previous works have leveraged metrics such as attention score contribution Zhang et al. (2024b) to select tokens to evict from the KV cache. One challenge with evicting tokens from the KV cache is that it irreversibly discards less important tokens - if these tokens become important later during generation, they cannot be recovered. KV cache merging Wang et al. (2024); Yuan et al. (2025) was proposed as a solution to compensate for the error induced by KV cache pruning. KV cache merging involves merging evicted KV cache entries (based on key similarity Wang et al. (2024) or historical attention scores Yuan et al. (2025)) instead of discarding them completely. In contrast with these works, our algorithm retains the full original KV cache (thereby allowing the model to refer back to any prior KV cache state exactly if they become important during generation), while approximating attention to less important clusters using the key centroids.
@@
Another common approach for leveraging sparsity for the KV cache is to retain the full KV cache and only load in the required KV cache entries for each decoding step. One prior work, QUEST Tang et al. (2024), grouped consecutive KV cache entries and derived representative vectors for the keys within each group, and compared the query with these representative vectors to decide which KV cache entries to load. Whereas QUEST clustered keys based on positional proximity, Squeezed Attention Hooper et al. (2024a) instead clustered keys based on semantic similarity, thereby allowing for precise identification of tokens which were likely to be high-scoring. Another related work, Tactic Zhu et al. (2025), performed clustering based on semantic similarity and then used distribution fitting to select critical KV cache entries. Other previous works have framed the query-key comparison as a retrieval problem in order to accerate attention using vector search methods Zhang et al. (2024a); Liu et al. (2024a); He et al. (2025). In contrast with existing methods, our approach aims to identify and load only a small subset of KV cache entries for exact attention computation, while approximating the attention to the remaining KV cache entries. Our algorithm clusters keys based on semantic similarity and then performs a fast centroid comparison to identify important keys, while also leveraging these representative centroids to approximate the attention to the keys which are not retrieved by the centroid lookup. A more detailed discussion of methods that sparsely load KV cache entries is provided in Appendix A.
@@ 3.1 Retrieving Important KV Cache Tokens
In order to be able to identify and retrieve important KV cache entries, we first cluster the keys based on semantic similarity using -means clustering. We then derive representative centroids for each cluster by taking the mean of all vectors in each cluster, which can then be used to quickly estimate the importance of each key cluster (as in Hooper et al. (2024a)). To identify whether the keys in cluster are important for a given query token , we compare the query with the key centroid for that cluster to estimate the attention score for the corresponding keys:
where is the key centroid for cluster and is the number of keys in cluster . We can then use the centroid attention score to identify the clusters which are likely to be high scoring, and only load in the corresponding keys from these clusters for exact attention computation. To decide which keys to retain, we sort the clusters based on the centroid scores, and then retain clusters until we hit a target token budget. Note that for models which support grouped query attention, we aggregate importance by averaging the estimated attention across query heads which share keys and values Ainslie et al. (2023).
One challenge when clustering the keys is the impact of the Rotary Positional Embeddings (RoPE), which rotates key vectors by different amounts depending on their position in the sequence. This means that semantically similar keys at different positions in the sequence may not be clustered together. To address this, we employ the Windowed RoPE strategy from He et al. (2025) when computing the centroids (which assumes a fixed relative positional difference between the query and the key vectors), thereby improving the clusterability of the key vectors. We then use a query vector rotated at a fixed position when performing the centroid lookup (as in He et al. (2025)).
@@ 3.3 Efficient Online Clustering
Additionally, to accelerate the update process, we sample additional initial centroids randomly from the appended tokens, and then leverage a fast initial assignment of the appended tokens to the nearest centroids and corresponding update of the centroids that these tokens map to (this is analogous to a batched version of the sequential -means algorithm MacQueen (1967)). This process only needs to be run for the newly appended tokens, which makes its runtime overhead minimal. After running this single-shot update, to ensure high clustering accuracy we run a small number of iterations of full -means over the final block to refine the centroids and centroid assignment.
@@ 3.4 Hierarchical Multipole Attention
To alleviate the overhead of the centroid lookup, while still providing an accurate approximation for the full KV cache, we extend the hierarchical centroid lookup method from Hooper et al. (2024a) to a hierarchical multipole approximation for the attention operation. We perform hierarchical -means clustering in order to derive progressively coarser grained cluster centroids. This hierarchical clustering is compatible with the fast online clustering update described in Section 3.3, as the blockwise clustering method can be applied at each level of the hierarchy. At each level of the centroid lookup, we identify a small number of promising key centroids (for which we need to compare with the next level of refined centroids), and then approximate the less important keys using the attention to the key centroids.
@@ 4 System Implementation
We implement custom Triton kernels for computing the centroid lookup, sparse Flash Decoding, and centroid replacement operations. These kernels follow the parallelization strategies from FlashAttention-2 Dao (2023), including splitting computation across attention heads and along the query sequence length dimension, and FlashDecoding Dao et al. (2023), where computation is split across the KV dimension and the intermediate results are reduced. We also build upon existing work on sparse FlashAttention kernels Pagliardini et al. (2023); Hooper et al. (2024a).
The first stage of our kernel implementation (the centroid lookup) is designed to compare the query with the key centroids in order to identify the important tokens (as in Hooper et al. (2024a)). After obtaining the attention scores to the centroids, we gather the centroid scores per-token and apply top-K to identify which top-scoring clusters of keys can be retained under a fixed token budget. After identifying the important and less important clusters, the next phase of our kernel implementation (sparse FlashDecoding) calculates the attention output from keys belonging to selected (important) clusters. This kernel accepts as input a tensor containing the indices of keys, and only loads the keys corresponding to these indices for exact attention computation.
@@ 5.1 Experimental Setup
We evaluate Multipole Attention on two long context reasoning datasets: LongBenchV2 Bai et al. (2024) and GSM-Infinite Zhou et al. (2025). LongBenchV2 contains complex real-world long-context questions across tasks such as document QA, long in-context learning, dialogue understanding, and code repository understanding. GSM-Infinite contains synthetic mathematical reasoning tasks that require multiple arithmetic operations. We leverage two open-source reasoning models for our evaluation: Qwen3-8B Yang et al. (2024) and DeepSeek-R1-Distil-Qwen-14B. We use YaRN scaling across our experiments to enable up to 128K context length Peng et al. (2023). We use the recommended decoding settings for both models, which is to use a temperature of 0.6 and a top-p value of 0.95, and to set the top-k value to 20 for Qwen3-8B. We report average results across 3 trials for all experiments.
We evaluate our approach using token budgets of 128 and 512. We fix the block size for clustering as tokens throughout our evaluation and set , and we set (meaning that we keep the last 128 to 256 tokens unclustered and we update the clusters every 128 decoding steps). We set the number of -means iterations as 10 for prefill and 3 for refinement with our fast cluster update, and we use random initialization for the centroids. We also leave the first 10 “Attention Sink” tokens untouched due to their disproportionate importance Xiao et al. (2023).
We compare our method with Squeezed Attention Hooper et al. (2024a) and QUEST Tang et al. (2024) as representative baselines for sparse attention. Note that for Squeezed Attention, we report results that leverage our optimized online clustering implementation in order to identify important KV cache entries among the generated output tokens (as well as windowed RoPE for improved key clustering), but without our method for approximating attention to clusters of tokens using the attention to the centroid. The token budgets for the baselines are adjusted to ensure that they retain the same number of tokens when factoring in the local buffer used by Multipole Attention. For Multipole Attention, we use one centroid per 16 tokens, and for QUEST we use a page size of 16. For Squeezed Attention, we use one centroid per 8 tokens for fair comparison, since it only has to store key centroids (whereas our method stores key and value centroids). For evaluation with the hierarchical variant of Multipole Attention, we use a two-level lookup with 1 centroid per 64 tokens for the first level, and with one centroid per 8 tokens for the second level (and when comparing with the hierarchical centroid lookup from Squeezed Attention, we run their method with twice as many centroids at each level for fair comparison with our approach).

### v2

@@ Abstract
Large Reasoning Models (LRMs) have shown promising accuracy improvements for complex problem-solving tasks. While these models have attained high accuracy by leveraging additional computation at test time, they need to generate long chain-of-thought reasoning in order to think before answering, which requires generating thousands of tokens. While sparse attention methods can help reduce the KV cache pressure induced by this long autoregressive reasoning, these methods can introduce errors which disrupt the reasoning process. Additionally, prior methods often pre-processed the input to make it easier to identify the important prompt tokens when computing attention during generation, and this pre-processing is challenging to perform online for newly generated reasoning tokens. Our work addresses these challenges by introducing Multipole Attention, which accelerates autoregressive reasoning by only computing exact attention for the most important tokens, while maintaining approximate representations for the remaining tokens. Our method first performs clustering to group together semantically similar key vectors, and then uses the cluster centroids both to identify important key vectors and to approximate the remaining key vectors in order to retain high accuracy. Additionally, we design a fast cluster update process to quickly re-cluster the input and previously generated tokens, thereby allowing for accelerating attention to the previous output tokens. We evaluate our method using emerging LRMs such as Qwen-8B and Deepseek-R1-Distil-Qwen2.5-14B, demonstrating that our approach can maintain accuracy on complex reasoning tasks even with aggressive attention sparsity settings. We also provide kernel implementations to demonstrate the practical efficiency gains from our method, achieving up to 4.5 speedup for attention in long-context reasoning applications. Our code is available at https://github.com/SqueezeAILab/MultipoleAttention.
@@ 1 Introduction
Test-time compute has emerged as a new dimension for scaling up Large Language Model (LLM) performance (snell2024scaling). By scaling the amount of computation used at test time, we can allow LLMs to think longer on harder problems in order to attain higher accuracy. Large Reasoning Models (LRMs), which are LLMs post-trained to have strong reasoning capabilities, typically reason by generating long chain-of-thought, where they produce a step-by-step chain to help guide themselves to the correct solution (wei2022chain; guo2025deepseek; qwq32b). This has been shown to provide substantial accuracy gains in domains such as math and coding which require problem solving capabilities (guo2025deepseek; qwq32b).
However, the long autoregressive generation required for reasoning also brings substantial inference efficiency challenges. Notably, the reasoning process leads to large Key/Value (KV) cache memory footprint (hooper2024kvquant). One common method to reduce the KV cache footprint is through sparse attention methods that only load important KV cache tokens at each generation step (tang2024quest; hooper2024squeezed). However, while these methods are promising for reducing the KV cache memory bottleneck, they lead to unacceptable accuracy loss. Prior work has also shown that problem solving tasks which require complex reasoning experience greater accuracy degradation when pruning the KV cache (liu2025can). Additionally, existing sparse attention methods often pre-process the KV cache from the prompt to make it easier to retrieve important prompt tokens during generation, and this pre-processing is challenging to perform online during generation for the previously generated reasoning tokens (hooper2024squeezed).
To address these challenges, we build on top of (hooper2024squeezed) and introduce Multipole Attention, which leverages sparse attention in order to only load the important KV cache tokens, while still approximating attention to the rest of the context to maintain high accuracy. Figure 1 provides an intuitive visualization of our algorithm, which first clusters the keys and computes a representative centroid for each cluster. When computing attention, our method compares the current query with the centroids to determine the importance of key tokens. For the highest scoring keys, we compute exact attention. For the remaining keys, we leverage the cluster centroid to approximate the attention score for all keys in that cluster, thereby retaining contextual information from the full KV cache. We also present a hierarchical generalization of our method which leverages progressively coarser grained centroids as we get further away from the query. To accelerate attention to the previously generated reasoning tokens, we also design a fast cluster update strategy that can be applied during generation, thereby allowing our method to accelerate attention to the previous output tokens.
@@
Multipole Approximation: Our algorithm clusters keys based on semantic similarity and represents all keys within a cluster with a representative key centroid (as in (hooper2024squeezed)). When computing attention, we then compare the current query with the key centroids to identify which keys are important for the current query; we only compute exact attention for these important keys, and we approximate the attention to the remaining keys using the attention scores with the cluster centroids. This allows our algorithm to maintain contextually relevant information from the full sequence, even with low KV cache budget. We also extend this to a hierarchical multipole approximation to improve the efficiency of the centroid comparison.
@@ 2.1 Reasoning Models
Reasoning models have emerged as a new paradigm for solving complex problem-solving tasks (guo2025deepseek; qwq32b). These models are trained to generate a chain-of-thought reasoning trajectory, which contains step-by-step reasoning to help guide the model to the correct answer (wei2022chain). The DeepSeek-R1 model (guo2025deepseek) pioneered a new reinforcement learning-based methodology for training reasoning models for problem solving tasks. Their model release also included smaller open-source models (distilled from their base model) based on the Llama3 (grattafiori2024llama) and Qwen2.5 model series (qwen2.5). There have also been several powerful reasoning models released with agentic tool-calling capabilites such as the QwQ-32B model (qwq32b), the Qwen3 model series (qwen2.5), and the Llama-Nemotron model series (bercovich2025llamanemotronefficientreasoningmodels).
@@ 2.2 Long Context Inference
Long-context length LLMs, which can support context lengths greater than 100K tokens, have seen widespread use for applications such as document question answering, summarization, literature review, and processing multi-turn conversations. There have been several closed-source models which support long context capabilities (achiam2023gpt; claude2), including Gemini (gemini15) which supports greater than 1 million context length. Newer open-source model families such as Llama3 (llama31) and Qwen2.5 (qwen2.5) also support greater than 100K context length. Another line of work has focused on extending the context windows of pretrained models beyond the original supported context length through scaling or adjusting the positional embeddings (chen2023extending; peng2023yarn).
@@ 2.3 KV Cache Compression
KV cache compression has emerged as a key algorithmic tool for enabling long context length inference. For long context inference scenarios, the KV cache becomes the main memory bottleneck, making KV cache compression crucial for efficient inference (hooper2024kvquant). Common approaches for compressing the KV cache include quantization (hooper2024kvquant; liu2024kivi), which aims to compress the representation for each token, and sparsification (tang2024quest; hooper2024squeezed; zhang2024h2o), which aims to only load important KV cache tokens.
@@
One previous approach for inducing sparsity to reduce KV cache size during inference is KV cache eviction, where less important tokens are evicted during the generation process to reduce memory consumption. Previous works have leveraged metrics such as attention score contribution (zhang2024h2o) to select tokens to evict from the KV cache. One challenge with evicting tokens from the KV cache is that it irreversibly discards less important tokens - if these tokens become important later during generation, they cannot be recovered. KV cache merging (wang2024model; yuan2025weightedkv) was proposed as a solution to compensate for the error induced by KV cache pruning. KV cache merging involves merging evicted KV cache entries (based on key similarity (wang2024model) or historical attention scores (yuan2025weightedkv)) instead of discarding them completely. In contrast with these works, our algorithm retains the full original KV cache (thereby allowing the model to refer back to any prior KV cache state exactly if they become important during generation), while approximating attention to less important clusters using the key centroids.
@@
Another common approach for leveraging sparsity for the KV cache is to retain the full KV cache and only load in the required KV cache entries for each decoding step. One prior work, QUEST (tang2024quest), grouped consecutive KV cache entries and derived representative vectors for the keys within each group, and compared the query with these representative vectors to decide which KV cache entries to load. Whereas QUEST clustered keys based on positional proximity, Squeezed Attention (hooper2024squeezed) instead clustered keys based on semantic similarity, thereby allowing for precise identification of tokens which were likely to be high-scoring. Another related work, Tactic (zhu2025tactic), performed clustering based on semantic similarity and then used distribution fitting to select critical KV cache entries. Other previous works have framed the query-key comparison as a retrieval problem in order to accelerate attention using vector search methods (zhang2024pqcache; liu2024retrievalattention; he20252). In contrast with existing methods, our approach aims to identify and load only a small subset of KV cache entries for exact attention computation, while approximating the attention to the remaining KV cache entries. Our algorithm clusters keys based on semantic similarity and then performs a fast centroid comparison to identify important keys, while also leveraging these representative centroids to approximate the attention to the keys which are not retrieved by the centroid lookup. A more detailed discussion of methods that sparsely load KV cache entries is provided in Appendix A.
@@ 3.1 Retrieving Important KV Cache Tokens
In order to be able to identify and retrieve important KV cache entries, we first cluster the keys based on semantic similarity using -means clustering. We then derive representative centroids for each cluster by taking the mean of all vectors in each cluster, which can then be used to quickly estimate the importance of each key cluster (as in (hooper2024squeezed)). To identify whether the keys in cluster are important for a given query token , we compare the query with the key centroid for that cluster to estimate the attention score for the corresponding keys:
where is the key centroid for cluster and is the number of keys in cluster . We can then use the centroid attention score to identify the clusters which are likely to be high scoring, and only load in the corresponding keys from these clusters for exact attention computation. To decide which keys to retain, we sort the clusters based on the centroid scores, and then retain clusters until we hit a target token budget. Note that for models which support grouped query attention, we aggregate importance by averaging the estimated attention across query heads which share keys and values (ainslie2023gqa).
One challenge when clustering the keys is the impact of the Rotary Positional Embeddings (RoPE), which rotates key vectors by different amounts depending on their position in the sequence. This means that semantically similar keys at different positions in the sequence may not be clustered together. To address this, we employ the Windowed RoPE strategy from (he20252) when computing the centroids (which assumes a fixed relative positional difference between the query and the key vectors), thereby improving the clusterability of the key vectors. We then use a query vector rotated at a fixed position when performing the centroid lookup (as in (he20252)).
@@ 3.3 Efficient Online Clustering
Additionally, to accelerate the update process, we sample additional initial centroids randomly from the appended tokens, and then leverage a fast initial assignment of the appended tokens to the nearest centroids and corresponding update of the centroids that these tokens map to (this is analogous to a batched version of the sequential -means algorithm (macqueen1967some)). This process only needs to be run for the newly appended tokens, which makes its runtime overhead minimal. After running this single-shot update, to ensure high clustering accuracy we run a small number of iterations of full -means over the final block to refine the centroids and centroid assignment.
@@ 3.4 Hierarchical Multipole Attention
To alleviate the overhead of the centroid lookup, while still providing an accurate approximation for the full KV cache, we extend the hierarchical centroid lookup method from (hooper2024squeezed) to a hierarchical multipole approximation for the attention operation. We perform hierarchical -means clustering in order to derive progressively coarser grained cluster centroids. This hierarchical clustering is compatible with the fast online clustering update described in Section 3.3, as the blockwise clustering method can be applied at each level of the hierarchy. At each level of the centroid lookup, we identify a small number of promising key centroids (for which we need to compare with the next level of refined centroids), and then approximate the less important keys using the attention to the key centroids.
@@ 4 System Implementation
We implement custom Triton kernels for computing the centroid lookup, sparse Flash Decoding, and centroid replacement operations. These kernels follow the parallelization strategies from FlashAttention-2 (dao2023flashattention), including splitting computation across attention heads and along the query sequence length dimension, and FlashDecoding (flashdecoding), where computation is split across the KV dimension and the intermediate results are reduced. We also build upon existing work on sparse FlashAttention kernels (pagliardini2023fast; hooper2024squeezed).
The first stage of our kernel implementation (the centroid lookup) is designed to compare the query with the key centroids in order to identify the important tokens (as in (hooper2024squeezed)). After obtaining the attention scores to the centroids, we gather the centroid scores per-token and apply top-K to identify which top-scoring clusters of keys can be retained under a fixed token budget. After identifying the important and less important clusters, the next phase of our kernel implementation (sparse FlashDecoding) calculates the attention output from keys belonging to selected (important) clusters. This kernel accepts as input a tensor containing the indices of keys, and only loads the keys corresponding to these indices for exact attention computation.
@@ 5.1 Experimental Setup
We evaluate Multipole Attention on two long context reasoning datasets: LongBenchV2 (bai2024longbench) and GSM-Infinite (zhou2025gsm). LongBenchV2 contains complex real-world long-context questions across tasks such as document QA, long in-context learning, dialogue understanding, and code repository understanding. GSM-Infinite contains synthetic mathematical reasoning tasks that require multiple arithmetic operations. We leverage two open-source reasoning models for our evaluation: Qwen3-8B (qwen2.5) and DeepSeek-R1-Distil-Qwen-14B. We use YaRN scaling across our experiments to enable up to 128K context length (peng2023yarn). We use the recommended decoding settings for both models, which is to use a temperature of 0.6 and a top-p value of 0.95, and to set the top-k value to 20 for Qwen3-8B. We report average results across 3 trials for all experiments.
We evaluate our approach using token budgets of 128 and 512. We fix the block size for clustering as tokens throughout our evaluation and set , and we set (meaning that we keep the last 128 to 256 tokens unclustered and we update the clusters every 128 decoding steps). We set the number of -means iterations as 10 for prefill and 3 for refinement with our fast cluster update, and we use random initialization for the centroids. We also leave the first 10 “Attention Sink” tokens untouched due to their disproportionate importance (xiao2023efficient).
We compare our method with Squeezed Attention (hooper2024squeezed) and QUEST (tang2024quest) as representative baselines for sparse attention. Note that for Squeezed Attention, we report results that leverage our optimized online clustering implementation in order to identify important KV cache entries among the generated output tokens (as well as windowed RoPE for improved key clustering), but without our method for approximating attention to clusters of tokens using the attention to the centroids. The token budgets for the baselines are adjusted to ensure that they retain the same number of tokens when factoring in the local buffer used by Multipole Attention. For Multipole Attention, we use one centroid per 16 tokens, and for QUEST we use a page size of 16. For Squeezed Attention, we use one centroid per 8 tokens for fair comparison, since it only has to store key centroids (whereas our method stores key and value centroids). For evaluation with the hierarchical variant of Multipole Attention, we use a two-level lookup with one centroid per 64 tokens for the first level, and with one centroid per 8 tokens for the second level (and when comparing with the hierarchical centroid lookup from Squeezed Attention, we run their method with twice as many centroids at each level for fair comparison with our approach).